OpenStudy (anonymous):

x−1/x =x What is x equal to? (hint: pair of conjugate complex solutions)

OpenStudy (freckles):

\[\frac{x-1}{x}=x ?\]

OpenStudy (freckles):

\[x \neq 0 \\ \frac{x-1}{x}=x \\ x \cdot \frac{x-1}{x} = x \cdot x \\ \cancel{x} \frac{x-1}{\cancel{x}}=x^2 \\ x-1=x^2 \] See if you can solve this quadratic